# STD-LLM: Understanding Both Spatial and Temporal Properties of Spatial-Temporal Data with LLMs

### v1

@@ Abstract
Spatial-temporal forecasting and imputation are important for real-world dynamic systems such as intelligent transportation, urban planning, and public health. Most existing methods are tailored for individual forecasting or imputation tasks but are not designed for both. Additionally, they are less effective for zero-shot and few-shot learning. While large language models (LLMs) have exhibited strong pattern recognition and reasoning abilities across various tasks, including few-shot and zero-shot learning, their development in understanding spatial-temporal data has been constrained by insufficient modeling of complex correlations such as the temporal correlations, spatial connectivity, non-pairwise and high-order spatial-temporal correlations within data. In this paper, we propose STD-LLM for understanding both spatial and temporal properties of Spatial-Temporal Data with LLMs, which is capable of implementing both spatial-temporal forecasting and imputation tasks. STD-LLM understands spatial-temporal correlations via explicitly designed spatial and temporal tokenizers as well as virtual nodes. Topology-aware node embeddings are designed for LLMs to comprehend and exploit the topology structure of data. Additionally, to capture the non-pairwise and higher-order correlations, we design a hypergraph learning module for LLMs, which can enhance the overall performance and improve efficiency. Extensive experiments demonstrate that STD-LLM exhibits strong performance and generalization capabilities across the forecasting and imputation tasks on various datasets. Moreover, STD-LLM achieves promising results on both few-shot and zero-shot learning tasks.
@@ 1 Introduction
Understanding both spatial and temporal properties of spatial-temporal data (e.g., the dependencies and changing patterns of data in spatial-temporal dimensions) is crucial for various real-world dynamic systems such as intelligent transportation [1], urban planning [2], and public health. In practice, spatial-temporal forecasting [3, 4] and imputation [5, 6] are the two most pivotal and common tasks relying on understanding the properties of spatial-temporal data. Specifically, precise spatial-temporal forecasting aids in effective traffic management and travel planning, and spatial-temporal imputation fills in missing data caused by unforeseen events during data collection and storage, enabling precise analysis of spatial-temporal patterns and supporting other dependent tasks. Although extensive studies have achieved satisfactory accuracy in spatial-temporal forecasting and imputation, they rely on extensive historical data for training. However, obtaining comprehensive datasets for all the studied regions is challenging due to the high cost of collecting and storing long-term data. Consequently, the zero-shot [7] and few-shot learning [8] capabilities are hindered for spatial-temporal forecasting methods across different regions. Moreover, existing methods are usually tailored to specific tasks and not designed for both forecasting and imputation. Each method requires domain expertise and task-specific designs, increasing costs and complicating the deployment of forecasting and imputation methods. To summarize, we need a method that possesses powerful zero-shot and few-shot learning capabilities and is versatile for both spatial-temporal forecasting and imputation tasks, making it practical and applicable in real-world scenarios.
@@
We have noticed that pre-trained Large Language Models (LLMs) [9, 10] are widely recognized for strong performance in zero-shot and few-shot learning in unfamiliar scenarios for various natural language processing tasks. Additionally, they are applicable across a diverse range of tasks [11, 12]. However, due to the significant differences in modality between spatial-temporal and textual data used in the pre-training process of LLMs, it is challenging to fine-tune LLMs to comprehend spatial-temporal data. In addition, training a spatial-temporal LLM from scratch is non-trivial due to the limited availability of spatial-temporal data compared to textual data. To address this issue, researchers have attempted to adapt pre-trained LLMs for understanding spatial-temporal data by creating tokenizers designed for fine-tuning LLMs [13, 14]. However, these methods offer insufficient representation for spatial-temporal data. Thus, they are less effective for understanding both spatial and temporal properties of spatial-temporal data with LLMs. First, existing methods, such as [13, 14], design tokens along the spatial dimension but ignore the tokens obtained from the temporal dimension. Second, topological connectivity information inherent in spatial-temporal data is less explored for empowering zero-shot learning capabilities. Thirdly, existing methods focus on modeling pairwise relations of tokens in the spatial dimension. They are less effective in capturing non-pairwise and higher-order spatial-temporal correlations. On the other hand, existing LLM-based spatial-temporal forecasting methods primarily concentrate on forecasting tasks and overlook the task of imputation. This limitation restricts their versatility and efficiency for practical applications.
@@
To solve the above limitations and challenges, we propose STD-LLM for understanding both spatial-temporal properties of Spatio-Temporal Data with LLMs. First, we create spatial-temporal embeddings for fully capturing temporal correlations and exploiting the topology structure of data along the spatial dimension. Second, we develop a spatial tokenizer and a temporal tokenizer to activate the ability of pre-trained LLM to understand spatial-temporal data from both the spatial and temporal dimensions while incorporating the topology structure of data. Lastly, we design a hypergraph learning module to capture non-pairwise and higher-order spatial-temporal correlations and reduce the computational cost, particularly in scenarios with large graph structures. Based on the explicitly designed tokenization and hypergraph learning equipped for fine-tuning the LLM, STD-LLM is not only versatile for spatial-temporal forecasting and imputation but also exhibits accurate few-shot and zero-shot learning capabilities on spatial-temporal data. In summary, our contributions can be summarized as follows:
@@
  • We propose STD-LLM for accurate spatial-temporal forecasting and imputation, as well as zero-shot and few-shot learning, by activating LLMs to understand both the spatial and temporal properties of ST data.
@@
  • We develop temporal tokenizers and spatial tokenizers to construct tokens from spatial and temporal perspectives to activate LLM to understand spatial-temporal data while incorporating the topology structure inherent in data.
  • We design a hypergraph learning module to model the non-pairwise and higher-order spatial-temporal correlations in the spatial-temporal data. By incorporating the hypergraph learning module, we not only enhance the overall performance but also improve efficiency.
@@ 2 Related Work
Spatial-Temporal Data Forecasting. As a problem with a wide range of applications, there is a lot of work for this problem. Initial studies[15, 16] utilizing conventional time-series analysis techniques which based on basic statistical features of time series. In recent years, with the rapid development of deep learning, a large number of models that can effectively model spatial-temporal dependencies have emerged. Several studies [17] address temporal dynamics using RNNs and their derivatives. To deal with spatial dependency, the data can be divided into grids [18, 19, 20, 21], and then CNNs are utilized to capture spatial correlations. This approach is somewhat effective, but not all data can be partitioned into grid form. In order to achieve a more general and effective model, researchers introduced the graph convolution model[22, 23] thereby implementing spatial feature aggregation based on adjacency matrices [24, 25, 26, 27, 28, 3]. AGCRN[3] uses individual weights and biases for each node during GCN feature transformations to learn patterns specific to each node, and to remove the constraints of a priori adjacency matrices while constructing dynamic graph networks, AGCRN also can learns adjacency matrices adaptively from data by node embedding. Models such as GMANZheng et al. [29],ASTGCN[30] and ASTGNN[4] utilize attention mechanisms to further deal with temporal and spatial correlations. PDFormer[31] devised a spatial self-attention module to model global semantic neighborhoods and geographically localized neighborhoods separately via graph masking, and proposed a delay-aware feature transformation module to explicit model the temporal delay of propagation. The current work has a fairly good forecasting accuracy, but most of the models do not take into account the few-shot learning, zero-shot learning.
@@
Spatial-Temporal Data Imputation. Initially, researchers used interpolation algorithms, KNN and other algorithms to accomplish the imputation of some simple scenes. In order to realize imputation in more complex scenes, successive works based on low-rank matrix completion[32, 33] and random forest[34] were proposed. LATC[35] effectively captures local and global trends in the data by combining low-rank matrix complementation with auto-regressive model. The development of deep learning has also contributed to the research on imputation. Brits[5] build an efficient self-supervised imputation model by capturing complex correlations between different features, time through bidirectional LSTM. E2GAN[36], GAIN[37]and other GAN[38] based models transforms the imputation problem into a generative problem. In addition, some VAE[39] based work recovers missing values by modeling the distribution of the data in the hidden space and recovering the hidden state obtained by sampling from the distribution. Such as mTAN[40] obtains a representation of the data in the hidden space through the attention between the reference point and the incomplete sequence. Recently the powerful effect of Diffusion Model[41] in the field of image generation attracts a large number of researchers, and some imputation work based on Diffusion Model has also appeared. CSDI[42] is based on the Conditional Diffusion Model, which uses the observable as a conditional input, and learns the correlation between time and space through the attention mechanism. Compared to CSDI, PriSTI[6] adds conditional feature extraction module and graph neural network, which enhances the utilization of conditional information and the construction of geographic prior knowledge. Existing imputation models face the same problem as forecasting models, both need to reduce the need for training samples and enhance generalization capabilities.
@@
Large Language Model. LLMs provide a powerful and unified framework for different tasks and have attracted a large number of researchers. There has been a significant amount of recent work validating the effectiveness of the LLM in other modalities. One Fits All(OFA)[12] demonstrated the feasibility of using large models for time series data through the experimental results of eight different temporal tasks, and put forward preliminary explanations. Articles such as Time-LLM[43] make more detailed modifications to LLM, using methods such as reprogramming, prompt pool, and contrast learning to align temporal and textual data. Models such as STLLM[14] and STGLLM[13] enhance LLM’s perception of spatial-temporal data by constructing tokens in the dimension of space, combined with the fine-tuning method of partial freezing. Existing spatial-temporal data large models have achieved certain results, confirming that LLM has the ability to handle spatial-temporal data. However, they underutilize spatial information and are difficult to accurately capture complex spatial dependencies, and there is much room for improvement.
@@ 3.1 Problem Definition
Forecasting task. Given the historical spatial-temporal data with $T$ time slices $\{\bm{X}_{t-T+1},\ldots,$ $\bm{X}_{t-1},$ $\bm{X}_{t}\}$ and the corresponding graph structure, the objective is to learn a function $f$ with parameters $\theta$ to forecast future spatial-temporal data $\{\bm{X}_{t+1},\ldots,\bm{X}_{t+P-2},\bm{X}_{t+P-1}\}$ for the next $P$ time slices.
$\{\bm{X}_{t+1},\ldots,\bm{X}_{t+P-2},\bm{X}_{t+P-1}\}=f_{\theta}(\{\bm{X}_{t-T+1},\ldots,\bm{X}_{t-1},\bm{X}_{t}\},\mathcal{G}).$ (1)
Imputation task. Given spatial-temporal data with $T$ time slices $\{\bm{\tilde{X}}_{t-T+1},\ldots,\bm{\tilde{X}}_{t-1,:},\bm{\tilde{X}}_{t,:}\}$ that contain missing values, the goal is to learn a function $f$ with parameters $\theta$ to estimate missing values.
[…]
@@ 3.2 Model Structure
As shown in Figure 1, our model focuses on accurate spatial-temporal forecasting and imputation via fine-tuning the LLM to understand the dependencies and evolving patterns in spatial-temporal data. To achieve this, we first develop spatial-temporal embedding to exploit the topology structure and periodicity of data. Based on this, we design temporal and spatial tokenizers to convert the spatial-temporal data into sequential tokens, enabling the LLM to learn the inherent spatial-temporal dependencies and evolving patterns of the data represented in the sequential token format. Lastly, we further incorporate a hypergraph learning module to effectively capture the non-pairwise and higher-order spatial-temporal correlations in data.
Spatial-Temporal Embedding. The spatial-temporal embedding consists of a time embedding and a topology-aware node embedding, which explores the static properties of data in both the spatial and temporal dimensions.
Time embedding: Two kinds of temporal information, i.e., time-of-day and day-of-week, are exploited for creating embedding dictionaries, $\mathbf{D}_{t}\in\mathbb{R}^{288\times d_{t}}$ and $\mathbf{D}_{w}\in\mathbb{R}^{7\times d_{t}}$ for time-of-day and day-of-week, where $d_{t}$ is the dimension of the time embedding. By the looking-up and concatenation along with broadcast operations, we finally obtain the time embedding $\mathbf{E}_{T}\in\mathbb{R}^{T\times N\times 2d_{t}}$.
topology-aware node embedding: It is designed to activate LLM to understand the topology structure in spatial-temporal data and facilitate LLM’s universality in handling data with different numbers of nodes.
The proposed topology-aware node embedding is obtained based on the eigendecomposition of the graph Laplacian matrix. There are two reasons for this design. First, as a matrix representation of the graph, the graph Laplacian matrix encapsulates the important information (i.e., connectivity, degree, and eigenvalues) about the graph, helping the model perceive the static spatial properties of data. Second, the graph Laplacian matrix has mutually orthogonal eigenvectors, helping the model modeling the connectivity of nodes.
Specifically, we define the Laplacian matrix $\mathbf{L}=\mathbf{I}-\mathbf{D}^{\frac{1}{2}}\mathbf{A}\mathbf{D}^{\frac{1}{2}}$, where $\mathbf{D}=\sum_{j=1}^{N}\mathbf{A}_{i,j}$ is the degree matrix and $\mathbf{I}$ is the identity matrix. The eigendecomposition of $\mathbf{L}$ yields $\mathbf{L}=\mathbf{V}\mathbf{\Lambda}\mathbf{V}^{-1}$, with $\mathbf{V}$ being the matrix of eigenvectors and $\mathbf{\Lambda}$ the diagonal matrix of eigenvalues. By selecting the eigenvectors corresponding to the top $K$ largest eigenvalues, we obtain $\mathbf{V}^{{}^{\prime}}\in\mathbb{R}^{N\times K}$. After passing through a linear layer, we broadcast to obtain the topology-aware node embedding $\mathbf{E}_{N}\in\mathbb{R}^{T\times N\times d_{n}}$:
@@
$\mathbf{V}^{{}^{\prime}}=\mathbf{V}[:,\mathbf{w}^{*}],\ \ \ \ \mathbf{E}_{N}=\mathbf{W}_{ne}\mathbf{V}^{{}^{\prime}}+\mathbf{b}_{ne},$

### v2

# STD-PLM: Understanding Both Spatial and Temporal Properties of Spatial-Temporal Data with PLM

@@ Abstract
Spatial-temporal forecasting and imputation are important for real-world intelligent systems. Most existing methods are tailored for individual forecasting or imputation tasks but are not designed for both. Additionally, they are less effective for zero-shot and few-shot learning. While pre-trained language model (PLM) have exhibited strong pattern recognition and reasoning abilities across various tasks, including few-shot and zero-shot learning, their applications in spatial-temporal data understanding has been constrained by insufficient modeling of complex correlations such as the temporal correlations, spatial connectivity, non-pairwise and high-order spatial-temporal correlations within data. In this paper, we propose STD-PLM for understanding both spatial and temporal properties of Spatial-Temporal Data with PLM, which is capable of implementing both spatial-temporal forecasting and imputation tasks. STD-PLM understands spatial-temporal correlations via explicitly designed spatial and temporal tokenizers. Topology-aware node embeddings are designed for PLM to comprehend and exploit the topology structure of data in inductive manner. Furthermore, to mitigate the efficiency issues introduced by the PLM, we design a sandglass attention module (SGA) combined with a specific constrained loss function, which significantly improves the model’s efficiency while ensuring performance. Extensive experiments demonstrate that STD-PLM exhibits competitive performance and generalization capabilities across the forecasting and imputation tasks on various datasets. Moreover, STD-PLM achieves promising results on both few-shot and zero-shot tasks.
@@ 1 Introduction
Understanding both spatial and temporal properties of spatial-temporal data is crucial for various real-world dynamic systems such as intelligent transportation  [1] and urban planning  [2]. In practice, spatial-temporal forecasting and imputation are the two most pivotal and common tasks. Specifically, precise spatial-temporal forecasting aids in effective traffic management and travel planning, while spatial-temporal imputation enables precise analysis of spatial-temporal patterns and supporting other dependent tasks. Although extensive studies have achieved satisfactory accuracy in spatial-temporal forecasting and imputation, they rely on extensive historical data for training. However, obtaining comprehensive datasets for all the studied regions is challenging due to the high cost of collecting long-term data. Consequently, the zero-shot  [3] and few-shot learning  [4] capabilities are hindered for spatial-temporal forecasting methods across different regions. Moreover, existing methods are usually tailored to specific tasks and not designed for both forecasting and imputation. Each method requires domain expertise and task-specific designs, increasing costs and complicating the deployment of forecasting and imputation methods. To summarize, we need a method that possesses powerful zero-shot and few-shot learning capabilities and is versatile for both spatial-temporal forecasting and imputation tasks, making it practical and applicable in real-world scenarios.
@@
We have noticed that PLM are widely recognized for strong performance in zero-shot and few-shot learning across a diverse range of tasks  [5]. However, due to the significant differences between spatial-temporal and textual data, it is challenging for PLM to comprehend spatial-temporal data. In addition, training a spatial-temporal PLM from scratch is non-trivial due to the limited availability of spatial-temporal data compared to textual data. To address this issue, researchers have attempted to adapt PLM for understanding spatial-temporal data  [6, 7]. However, these methods offer insufficient representation for spatial-temporal data. Thus, they are less effective for understanding both spatial and temporal properties of spatial-temporal data with PLM. First, existing PLM-based methods design tokens along the spatial dimension but ignore the tokens obtained from the temporal dimension. Second, topological connectivity information inherent in spatial-temporal data is less explored. Thirdly, existing methods do not take into account that under the large embedding dimensions of the PLM, as the number of tokens increases, the training and inference costs of the model will increase rapidly. On the other hand, existing PLM-based spatial-temporal forecasting methods primarily concentrate on forecasting tasks and overlook the task of imputation. This limitation restricts their versatility and efficiency for practical applications.
@@
To solve the above limitations and challenges, we propose STD-PLM for understanding both spatial-temporal properties of Spatio-Temporal Data with PLM. First, we create spatial-temporal embeddings for fully capturing temporal correlations and exploiting the topology structure of data. Second, we develop a spatial tokenizer and a temporal tokenizer to activate the ability of pre-trained PLM to understand spatial-temporal data from both the spatial and temporal dimensions while incorporating the topology structure of data. Lastly, we design the sandglass attention module with a constrained loss function, which can effectively model more important and sparser non-pairwise and higher-order correlations, significantly reducing computational costs. Based on the explicitly design, STD-PLM is not only versatile for spatial-temporal forecasting and imputation but also exhibits accurate few-shot and zero-shot learning capabilities on spatial-temporal data. In summary, our contributions can be summarized as follows:
@@
  • We propose STD-PLM for accurate spatial-temporal forecasting and imputation, as well as zero-shot and few-shot learning, by activating PLM to understand both the spatial and temporal properties of spatial-temporal data.
@@
  • We develop spatial-temporal tokenizers to construct tokens from spatial and temporal dimensions while incorporating the topology structure, thereby enabling the PLM to perceive not only fine-grained spatial information but also coarse-grained overall temporal information.
  • We design the sandglass attention module and its constrained loss function, which significantly reduces computational overhead while ensuring performance.
@@ 2 Related Work
Spatial-Temporal Data Forecasting. In recent years, a large number of models that can effectively model spatial-temporal dependencies have emerged. Several studies [8] address temporal dynamics using RNNs. To deal with spatial dependency, the data can be divided into grids [9, 10, 11, 12], and then CNNs are utilized to capture spatial correlations. But not all data can be partitioned into grid form. In order to achieve a more general and effective model, researchers introduced the graph convolution model [13, 14] thereby implementing spatial feature aggregation based on adjacency matrices [15, 16, 17, 18, 19, 20]. Models such as GMAN [21],ASTGCN [22] and ASTGNN [23] utilize attention mechanisms to further deal with temporal and spatial correlations. PDFormer [24] proposed a delay-aware feature transformation module to explicit model the temporal delay of propagation. The current work has a fairly good forecasting accuracy, but most of the models do not take into account the few-shot learning, zero-shot learning.
@@
Spatial-Temporal Data Imputation. Initially, in order to realize imputation in more complex scenes, successive works based on low-rank matrix completion [25, 26] were proposed. LATC [27] effectively captures local and global trends in the data by combining low-rank matrix complementation with auto-regressive model. The development of deep learning has also contributed to the research on imputation. Brits [28] build an efficient self-supervised imputation model through bidirectional LSTM. E2GAN [29], GAIN [30]and other GAN [31] based models transforms the imputation problem into a generative problem. In addition, some VAE [32] based works [33] recovers missing values by modeling the distribution of the data in the hidden space. Recently the powerful effect of Diffusion Model [34] in the field of image generation attracts a large number of researchers. CSDI [35] is based on the Diffusion Model, which recovery of missing values from noise by a well-designed denoising network. Compared to CSDI, another diffusion model PriSTI [36] enhances the utilization of conditional information and the construction of geographic prior knowledge. Existing imputation models face the same problem as forecasting models, both need to reduce the need for training samples and enhance generalization capabilities.
@@
Pre-trained Language Model. PLM provide a powerful and unified framework for different tasks. There has been a significant amount of recent work validating the effectiveness of the PLM in other modalities. One Fits All(OFA) [37] demonstrated the feasibility of using PLM for time series data through the experimental results of eight different temporal tasks, and put forward preliminary explanations. Articles such as Time-LLM [38, 39] make more detailed modifications to PLM, using methods such as reprogramming, prompt pool, and contrast learning to align temporal and textual data. Models such as STLLM [7] and STGLLM [6] enhance PLM’s perception of spatial-temporal data by constructing tokens in the dimension of space. Existing PLM-based models have achieved certain results, confirming that PLM has the ability to handle spatial-temporal data. However, they underutilize spatial information and are difficult to accurately capture complex spatial dependencies, and there is much room for improvement.
Sandglass Attention. There are some prior works whose methods are similar to the sandglass attention, such as SSTBAN [40] and CrossFormer [41], which both learn higher-order correlations through learnable reference points combined with an attention mechanism to improve model efficiency. Although these studies have achieved good results, they still have some limitations, such as not fully utilizing the adjacency relationships in the original graph and being unable to achieve zero-shot learning between different graphs.
@@ 3.1 Problem Definition
Spatial-Temporal data. Considering spatial-temporal data with $T$ time slices and $N$ nodes, we represent it as $\bm{\mathcal{X}}=\{\bm{X}_{1},\bm{X}_{2},\ldots,\bm{X}_{T}\}\in\mathbb{R}^{T\times N\times C}$. Here, $\bm{X}_{t}=\{\bm{x}_{t,1},\bm{x}_{t,2},\ldots,\bm{x}_{t,N}\}\in\mathbb{R}^{N\times C}$, where $\bm{x}_{t,n}$ represents the feature of node $n$ at time $t$, and $C$ is the number of features.
@@
Forecasting task. Given the historical spatial-temporal data with $T$ time slices $\bm{\mathcal{X}}$ and the corresponding graph structure, the objective is to forecast future data $\{\bm{X}_{t+1},\ldots,\bm{X}_{t+T-2},\bm{X}_{t+T-1}\}$ for the next $T$ time slices.
Imputation task. Given incomplete spatial-temporal data with $T$ time slices $\bm{\mathcal{X}}$, mask tensor $\bm{\mathcal{M}}$ and the corresponding graph structure, the goal is to estimate complete data $\{\bm{\hat{X}}_{t-T+1},\ldots,$ $\bm{\hat{X}}_{t-1},$ $\bm{\hat{X}}_{t}\}$.
Few-shot and Zero-shot task. The term ”few-shot" refers to training a model based on a limited amount of data. The term "zero-shot" denotes the direct testing of a model trained on a source dataset on a destination dataset, without any additional training.
@@ 3.2 Model Structure
As shown in Figure 1, our model is tailored for accurate spatial-temporal forecasting and imputation by fine-tuning PLM to understand the dependencies and evolving patterns in spatial-temporal data. To achieve this, we first develop a inductive spatial-temporal embedding module to exploit the topology structure and periodicity of data. Based on this, we design temporal and spatial tokenizers to convert the spatial-temporal data into sequential tokens, activating the ability of PLM to understand the inherent spatial, temporal and spatial-temporal correlations of the data represented in the sequential token format. Lastly, we further incorporate sandglass attention module consisting of a precoder and a decoder to not only improve the model efficiency, but also further capture the non-pairwise and higher-order spatial-temporal correlations.
3.2.1 Spatial-Temporal Embedding.
Spatial-temporal data exhibit heterogeneity at different nodes and time. Thus, we design spatial-temporal embeddings, comprising periodic-aware time embedding and topology-aware node embedding, to represent nodes and time in an inductive manner.
Topology-aware node embedding: In the design of node embeddings, we need them not only to reflect the static characteristics of each node but also to incorporate the topology structure to express the relationships between nodes. At the same time, the node embeddings must also possess inductive learning capabilities across different graph structures.
Noticing that the graph Laplacian matrix has two properties that align well with our needs, first, the graph Laplacian matrix contains important information about the graph’s structure, such as degree and connectivity. Second, the eigenvectors of the Laplacian matrix are orthogonal, which can effectively distinguish different nodes. An initial idea is to use the eigenvectors of the Laplacian matrix as embeddings for each node. However, considering that if the dimensionality of the embeddings is related to the graph’s order $N$, it would not be transferable across different graphs. Therefore, we choose to generate node embeddings based on the eigenvectors corresponding to the $K$ largest eigenvalues.
Specifically, we define the normalized Laplacian matrix $\mathbf{L}=\mathbf{I}-\mathbf{D}^{\frac{1}{2}}\mathbf{A}\mathbf{D}^{\frac{1}{2}}$, where $\mathbf{D}=\sum_{j=1}^{N}\mathbf{A}_{i,j}$ is the degree matrix and $\mathbf{I}$ is the identity matrix. The eigendecomposition of $\mathbf{L}$ yields $\mathbf{L}=\mathbf{V}\mathbf{\Lambda}\mathbf{V}^{-1}$, with $\mathbf{V}$ being the matrix of eigenvectors and $\mathbf{\Lambda}$ the diagonal matrix of eigenvalues. By selecting the eigenvectors corresponding to the top $K$ largest eigenvalues, we obtain $\mathbf{V}^{{}^{\prime}}\in\mathbb{R}^{N\times K}$. After passing through a linear layer, we broadcast to obtain the topology-aware node embedding $\mathbf{E}_{N}\in\mathbb{R}^{T\times N\times d_{n}}$:
@@
$\mathbf{V}^{{}^{\prime}}=\mathbf{V}[:,\mathbf{w}^{*}],\ \ \ \ \mathbf{E}_{N}=\mathbf{W}_{ne}\mathbf{V}^{{}^{\prime}}+\mathbf{b}_{ne},$ (2)